# ChemBOMAS: Accelerated BO in Chemistry with LLM-Enhanced Multi-Agent System

### v1

@@ Appendix D Wet Experiment Detail Protocol
Appendix E Additional Results Analysis
E.1 More Benchmark Results
As described in the main text, for comparison with baseline methods(here, traditional BO, LA-MCTS, LLAMBO), we conducted evaluations on four datasets. The configuration of baseline methods is detailed in Section F of the Supplementary Material. The results are shown in the Figure 4. Given the relatively small search space (see Table 1), even random selection may sample reasonably good variable combinations after numerous iterations. Consequently, we constricted our experiments to five rounds to evaluate the initialization performance of each algorithm during the early optimization stage.
E.2 Additional Results of Data Modules’s Prediction Model
This study evaluated the effectiveness of a prediction model (the Data Module)-based BO initialization strategy through experiments on the Suzuki reaction dataset (see Table 1). The experiment dataset comprised 15 types of Suzuki reactions within distinct reactant combinations, but all reaction types shared a variable space for reaction conditions. The Suzuki reaction dataset was partitioned into training and test datasets according to the types of reactant combinations: four types of reactant combinations were used to train an auxiliary predictive model to guide BO initialization, while the remaining 11 types of reactions constituted the test set to assess the performance of BO with different initialization strategies. The dataset division strategy was designed to mimic real-world scenarios, specifically predictions for previously unseen reactant combinations. Figure 5 illustrates the comparative initialization performance of our method and random sampling. The experimental results indicate that although the auxiliary predictive model exhibited a relatively low Root Mean Square Error (RMSE), the BO initialized by the predictive model captured the candidate points with high target values across different Suzuki reaction types in most test instances. Furthermore, compared to random sampling initialization (indicated by green lines in the figure), the proposed ChemBOMAS initialization strategy (indicated by red crosses in the figure) consistently provided superior initial candidate points for BO.
Appendix F Configuration of Baseline Methods
Traditional BO We implemented a traditional BO baseline with the BoTorch framework, using a Gaussian Process (GP) surrogate model with a Matérn kernel, a constant prior, and a qLogEI acquisition function ( = 0.001). The input of BO, namely the reaction condition variables, was encoded with simple one-hot vectors; we did not employ more elaborate encodings because previous studies report only marginal benefits Taylor et al. [2023], Shields et al. [2021]. Furthermore, the batch size of each iteration, including the initial round, was fixed at 5, and the initial population was generated by random sampling.
LLAMBO The LLAMBO implementation was based on the default configurations provided in the original authors’ publicly available codebase Liu et al. [2024]. A notable deviation from the original study was necessitated by the deprecation of the LLM API used therein. Consequently, we integrated a more recent and advanced LLM API, specifically GPT-4O OpenAI et al. [2024], for all LLM-dependent operations within the LLAMBO framework.
LA-MCTS For LA-MCTS, we utilized the open-source code from the original publication Wang et al. [2020], with hyperparameters adapted for optimal performance on chemical datasets. The algorithm was initialized with 5 randomly selected samples. The Support Vector Machine (SVM) component within LA-MCTS employed a Radial Basis Function (RBF) kernel. The splitting threshold for the MCTS was adjusted according to the dataset size, specifically set to 50 for larger datasets and 20 for smaller ones, to ensure appropriate tree expansion. The exploration constant, , was configured to 10. Sampling was performed using the TuRBO algorithm Eriksson et al. [2019], with its internal hyperparameters retained at their default values as specified in its original implementation.
Appendix G Detail of the Prompts
As outlined in the main text, our methodology leverages large language models to support several critical tasks in reaction optimization, such as analyzing literature, assessing parameter significance, and understanding physicochemical properties to inform the construction of a hierarchical optimization tree. This appendix section presents a detailed overview of the specific prompts designed to guide the LLM in executing these crucial Tasks.
G.1 Prompts of Knowledge Module
The Knowledge Module, as described in Section A of the Supplementary Material, employs the LLM to systematically analyze chemical literature and physicochemical data. This involves ranking the impact of various reaction parameters and classifying components based on their physicochemical properties.
Impact Sort Prompt The Impact Sort prompt is designed to instruct the LLM to evaluate and rank the relative importance of different reaction parameters (e.g., reactant, catalyst, ligand, solvent, base) on the outcome of a specific target reaction. The parameters ranking is derived from the LLM’s synthesis of information from relevant chemical literature and its general chemical knowledge. An example of the prompt used to elicit the parameter significance ranking is shown below: Prompt for Parameter Impact Ranking Based on Provided Literature. Objective:
Analyze the provided scientific literature excerpts related to the Suzuki-Miyaura cross-coupling reaction. Your task is to rank the following reaction parameters in descending order of their reported impact or significance on the yield.
Crucial Instruction: Your analysis, conclusions, and the resulting ranking MUST be based solely and exclusively on the information presented within the provided literature excerpts below. Do not use any external knowledge, pre-existing understanding, or general chemical intuition you might possess. Focus strictly on what the given texts state or imply regarding the influence of each parameter.
Reaction Parameters to Rank:
[Type of Solvent]
[Type of Base]
Provided Literature:
[LITERATURE_1]
[LITERATURE_2]
While toolkits exist that allow LLMs to actively retrieve literature from sources such as arXiv and Google Scholar, thereby enabling automated construction of a vector library for retrieval-augmented generation (RAG), these automated retrieval approaches face significant limitations. Many publishers impose paywalls or lack accessible APIs, and automated retrieval poses potential risks of data leakage during the evaluation process. Therefore, we opted to manually curate and provide the literature to ensure data integrity and rationality.
Variable Candidates Classification Prompt The prompt guides the LLM to identify key physicochemical properties of each variable and cluster variable candidates based on their similarity in the physicochemical properties. Below is an example of the prompt for variable candidates classification.
Prompt for Variable Candidates Classification Based on Physicochemical Data Objective:
Classify the provided list of candidate chemical substances according to the [Specified_physicochemical_Properties]. Your primary method for classification must be the utilization of quantitative data that would typically be found in a comprehensive physicochemical property database.
Crucial Instructions:
Prioritize Quantitative Data: For each substance and property, you should first attempt to classify it based on specific, measurable, quantitative values (e.g., pKa for basicity/acidity, dielectric constant for polarity, boiling point for volatility, specific functional group counts).
Minimize General Knowledge/Intuition: Avoid relying on your general, unquantified chemical knowledge or intuition. If a quantitative value from the "database" directly supports a classification, state that. If a direct value isn’t typically used for a category but strong structural indicators (which could be quantified, e.g., number of H-bond donors) point to it, explain this as an inference based on data-like principles.
Adhere to Provided Categories: Classify substances strictly into the categories provided for each property. If a substance doesn’t clearly fit or straddles categories based on (assumed) data, note this ambiguity.
Candidate Substances to Classify:
[CANDIDATE_SUBSTANCES_LIST]
Available Tools:
[GoogleScholarToolkit], [ArxivToolkit], [PubMedToolkit], [PubChemToolkit], [WikipediaToolkit]
G.2 Prompt of Data Module
As detailed in the main text, our pre-training phase employs a conditional prediction task. Given the reactants and products, the model’s objective is to predict the corresponding reaction conditions. This process utilizes a Causal Language Modeling (CLM) loss, where the model learns to predict the next token in the sequence of reaction conditions.
To provide concrete examples of the input format for this task, this appendix section (Section G.2) presents a selection of prompts utilized during the pre-training phase. These prompts typically consist of the reactants, products, and the target reaction condition sequence that the model is trained to predict. Furthermore, in line with the methodology described in the main text, these input sequences are augmented with functional group annotations (generated via RDKit) to enhance the model’s chemical awareness; the augmentation of the prompt is also reflected in the examples provided below.
Prompt of Condition Prediction Pretraining
To further clarify the input structure for this prediction task, the following examples demonstrate the format used: title=Condition Prediction Pre-training Prompts "reaction": "Here is a chemical reaction.
Reactants are: C1=CC=CC=2C3=CC=CC=C3N(C12)CC#C,BrC#CCCCCO.
Product is: C1=CC=CC=2C3=CC=CC=C3N(C12)CC#CC#CCCCCO.
Reaction type is Cadiot-Chodkiewicz coupling.",
"condition": "The reaction conditions of this reaction are:
Solvent: O,CN(C=O)C,CN(C=O)C. Catalyst: Cl[Cu]. Atmosphere: N#N. Additive: C(C)N,[Na]Cl,Cl.NO. ", "reaction_type": "Cadiot-Chodkiewicz coupling",
Prompt of Functional Groups Information
We utilize RDKit to extract functional group information from the SMILES representation of the reactants. This information includes the type of functional groups present in each reagent and their transformations during the reaction. By incorporating this data, we enhance the input to the LLM, providing it with richer chemical context.
Below is an example of the functional group information from the reactants in the Suzuki coupling dataset to augment the prompt. title=An Example of functional group information
Functional groups information:
Reactant: [CCc1cccc(CC)c1.Clc1ccc2ncccc2c1]. Functional Group: Chlorine([R]Cl). Count: 1. Reactant: [Cc1ccc2c(cnn2C2CCCCO2)c1B(O)O]. Functional Group: Alcohol([R]O). Count: 2. Reactant: [Cc1ccc2c(cnn2C2CCCCO2)c1B(O)O]. Functional Group: Ether([R]O[R]). Count: 1. Reactant: [Cc1ccc2c(cnn2C2CCCCO2)c1B(O)O]. Functional Group: BoricAcid([R]B(O)O). Count: 1.
Product: [Cc1ccc2c(cnn2C2CCCCO2)c1-c1ccc2ncccc2c1]. Functional Group: Alcohol([R]O). Count: 2. Product: [Cc1ccc2c(cnn2C2CCCCO2)c1-c1ccc2ncccc2c1]. Functional Group: Ether([R]O[R]). Count: 1. New Functional Groups: None. Lost Functional Groups: Functional Group: Chlorine([R]Cl). Count: 1. Lost: 1. Functional Group: BoricAcid([R]B(O)O). Count: 1. Lost: 1.
Appendix H Example of Knowledge Module Results
H.1 Example of Impact Sort Result
This section presents an "Impact Sort" result generated by an LLM. The ranking task given to the LLM was to analyze provided scientific literature excerpts concerning the Suzuki-Miyaura cross-coupling reaction. Based solely on these excerpts, the LLM was asked to rank reaction variables in descending order of their reported influence or significance on the reaction yield.
The LLM’s response, detailing the ranking and justifications, is as follows: Ranking of Reaction Variables by Impact on Yield (Most to Least Impactful):
1.
Catalyst
2.
Ligand
3.
Base
4.
Solvent
5.
Temperature
Justification Based on Provided Literature Excerpts:
1.
Ligand: The provided texts place enormous emphasis on the role of the ligand in determining reaction success and yield. …
2.
Catalyst: The catalyst, typically a palladium species, is fundamental to the reaction. Its nature, concentration, and activity are repeatedly linked to the reaction outcome. …
H.2 Example of Variable Candidates Classification
This section provides an example of the classification of various base candidates conducted by an LLM. The clustering task was to classify a list of candidate bases according to their pH in solution, based on provided literature information and pKa values.
The LLM’s response, detailing the classification and justifications, is as follows:
Summary of Classifications based on pH:
1.
Basic
2.
t-BuOK: Strongly Basic
3.
: Basic
4.
KOAc: Basic
5.
CsOAc: Basic
6.
NaOPh: Basic
7.
NaOTf: Neutral
Justification of some substance
(Sodium methoxide): The conjugate acid of methoxide () is methanol (). The pKa of methanol is approximately 15.5. Since methanol is a very weak acid, its conjugate base, methoxide, is a strong base. Therefore, in solution will be Basic.
t-BuOK (Potassium tert-butoxide): The conjugate acid of tert-butoxide () is tert-butanol (t-BuOH). The pKa of tert-butanol is approximately 16.54 to 19. As tert-butanol is a very weak acid, its conjugate base, tert-butoxide, is a very strong base. Thus, t-BuOK in solution will be Strongly Basic.
H.3 Example of Search Space Decomposition Result
Here, we present an example of the search space decomposition on the Buchwald-Hartwig coupling dataset. This search space decomposition is generated by the Knowledge Module. As specifically emphasized in the prompt, its generation process relies entirely on knowledge sourced from the literature. For brevity, only a portion of the decomposition is displayed. An example of decomposition "Buchwald": "Solvent": "CS(C)=O": "iupac": "methylsulfinylmethane", "cid": 679, "subset": 0 , "Base": "CN1CCCN2CCCN=C12": "iupac": "1-methyl-2,3,4,6,7,8-hexahydropyrimido[1,2-a]pyrimidine", "cid": 123583, "subset": 0 , "CN(C)C(=NC(C)(C)C)N(C)C": "iupac": "2-tert-butyl-1,1,3,3-tetramethylguanidine", "cid": 3571581, "subset": 0 , "CCN=P(N=P(N(C)C)(N(C)C)N(C)C)(N(C)C)N(C)C": "iupac": "N-[dimethylamino-ethylimino-[[tris(dimethylamino)-lambda5-phosphanylidene]amino]-lambda5-phosphanyl]-N-methylmethanamine", "cid": 3393106, "subset": 1 , "Ligand": "CC(C)c1cc(C(C)C)c(-c2ccccc2P(C(C)(C)C)C(C)(C)C)c(C(C)C)c1": "iupac": "ditert-butyl-[2-[2,4,6-tri(propan-2-yl)phenyl]phenyl]phosphane", "cid": 11618717, "subset": 0 , "CC(C)c1cc(C(C)C)c(-c2ccccc2P(C2CCCCC2)C2CCCCC2)c(C(C)C)c1": "iupac": "dicyclohexyl-[2-[2,4,6-tri(propan-2-yl)phenyl]phenyl]phosphane", "cid": 11155794, "subset": 1 , "COc1ccc(OC)c(P([C@]23C[C@H]4C[C@H](C[C@H](C4)C2)C3)[C@]23C[C@H]4C[C@H](C[C@H](C4)C2)C3)c1-c1c(C(C)C)cc(C(C)C)cc1C(C)C": "iupac": "bis(1-adamantyl)-[3,6-dimethoxy-2-[2,4,6-tri(propan-2-yl)phenyl]phenyl]phosphane", "cid": 60144828, "subset": 0 , "COc1ccc(OC)c(P(C(C)(C)C)C(C)(C)C)c1-c1c(C(C)C)cc(C(C)C)cc1C(C)C": "iupac": "ditert-butyl-[3,6-dimethoxy-2-[2,4,6-tri(propan-2-yl)phenyl]phenyl]phosphane", "cid": 44233348, "subset": 0

### v2

@@ H.2 Wet Experiment Detail Protocol
Appendix I Additional Results Analysis
I.1 Impact of Batch Size
To investigate the influence of the number of samples per iteration on the performance of ChemBOMAS, we conducted an ablation study by varying the batch size. Specifically, we configured the batch sizes to be 0.05%, 0.1%, 0.2%, and 0.4% of the total dataset volume for each reaction. It is important to note that for the Buchwald and Buchwald datasets, which have a significantly smaller number of data points, a batch percentage of 0.05% resulted in a batch size of less than 1. Consequently, experiments for this specific setting were omitted for these two datasets.
Dataset Batch % Batch Size Best Found Initial Value Time Iter. of Best Suzuki 0.05 3 96.15 72.98 105.18 3 0.10 5 96.15 72.98 158.03 3 0.20 10 96.15 72.98 318.11 3 0.40 20 96.15 72.98 646.02 3 Arylaton 0.05 2 80.67 78.71 59.24 37 0.10 3 80.64 78.71 95.01 39 0.20 6 81.65 78.71 240.63 39 0.40 12 81.25 78.71 755.86 29 Buchwald 0.05 - - - - - 0.10 1 79.58 75.55 10.51 25 0.20 2 79.39 75.55 24.43 23 0.40 4 79.60 75.55 56.30 25 Buchwald 0.05 - - - - - 0.10 1 56.81 53.34 8.77 2 0.20 2 56.81 53.34 20.57 2 0.40 4 56.81 53.34 49.02 2
The experimental results, as detailed in Table 6, revealed that while there was a marginal improvement in the best-found values with an increasing batch size, the overall impact on optimization performance was minimal. This observation underscored a critical strength of ChemBOMAS: its ability to efficiently navigate the variable space and identify optimal or near-optimal parameter combinations using a remarkably small subset of data in each iteration.
Conversely, a clear trend emerged regarding computational cost: the runtime increased substantially with larger batch sizes. Given the negligible gains in the optimal solution found, a larger batch size presented an unfavorable trade-off. Therefore, to balance computational efficiency and performance, we selected a batch percentage of 0.1% for our main experiments, as it demonstrated the capability of ChemBOMAS to achieve excellent results with minimal data sampling per round.
I.2 Impact of Prior Data Volume
To ascertain the influence of prior data volume on the fine-tuning process, we conducted a comprehensive ablation study. We fine-tuned the pre-trained Large Language Model on five distinct proportions of the four datasets: 0.25%, 0.5%, 1.0%, 2.0%, and 4.0%. The model’s predictive performance was evaluated with MSE, MAE, and as key metrics. Here, to more clearly demonstrate the impact of pseudo-data quality on Bayesian optimization, we have removed the knowledge-driven module.
Prior Data Suzuki Arylation Buchwald MSE MAE R MSE MAE R MSE MAE R 0.25% 1205.19 27.80 -0.53 793.92 24.20 -0.07 737.64 23.50 0.01 0.5% 774.70 21.13 0.02 1016.37 26.53 -0.36 617.28 19.81 0.17 1.0% 633.68 19.47 0.20 650.00 19.55 0.13 593.76 18.52 0.20 2.0% 479.09 15.92 0.39 462.52 15.75 0.38 365.60 13.98 0.51 4.0% 360.02 13.44 0.54 286.56 11.97 0.62 248.44 11.28 0.67
The empirical results, as presented in Table 7, demonstrate a clear and positive correlation between the volume of fine-tuning data and the model’s predictive accuracy. Specifically, we observed a monotonic improvement across all metrics as the data percentage increased. A critical threshold was identified at the 1.0% data level. At this point, the values for both the Suzuki and Arylation datasets became positive for the first time, signifying that the model’s predictions had surpassed the explanatory power of a simple mean-based model. Given that the scores for all three datasets converged to a reasonable performance level (approximately 0.2) at this stage, we concluded that 1.0% represents an efficient and effective baseline for prior data quantity, balancing model fidelity with data utilization. Therefore, this level was adopted for subsequent experiments.
We further investigated how the volume of the fine-tuned model, when leveraged as a surrogate for generating pseudo-data, affects the performance of a downstream Bayesian optimization task. For this, the models trained with 1%, 2%, and 4% of the prior data were utilized to guide the optimization process on four distinct reaction datasets: Suzuki, Arylation, Buchwald, and Buchwald.
Dataset Prior Data (%) Volume Size Best Found Initial Value Iteration of Best Suzuki 1.0 50 88.98 65.95 37 2.0 100 88.99 70.25 36 4.0 200 91.41 52.13 40 Arylation 1.0 30 79.67 45.98 40 2.0 60 81.65 48.95 40 4.0 120 83.81 37.11 31 Buchwald 1.0 7 79.63 54.07 31 2.0 14 79.56 41.14 39 4.0 28 80.91 32.36 15 Buchwald 1.0 7 56.81 12.87 11 2.0 14 53.95 15.59 38 4.0 28 53.93 16.22 40
The outcomes, summarized in Table 8, reveal a nuanced and somewhat counter-intuitive trend. While increasing the fine-tuning data from 1% to 4% (e.g., from 50 to 200 samples for Suzuki) does yield modest improvements in the "Best Found" values for Suzuki and Arylation, the enhancement is not proportional to the four-fold increase in data. More strikingly, for the Buchwald and Buchwald datasets, this trend does not hold. For instance, on Buchwald, the model fine-tuned with 1% of the data achieved a superior result (56.81) compared to the models trained on 2% (53.95) and 4% (53.93) of the data. Similarly, for Buchwald, the 1% model’s performance (79.63) was marginally better than the 2% model (79.56).
These findings suggest that simply increasing the volume of data for training the surrogate model does not guarantee enhanced performance in Bayesian optimization. Beyond a certain point, it appears to yield diminishing returns and can even be detrimental to the optimization outcome. This may indicate a complex interplay between the surrogate model’s accuracy and its ability to generalize across the search space, suggesting that a more moderately-sized, yet sufficiently representative, prior dataset may be optimal for guiding the exploration-exploitation balance in our optimization framework.
Appendix J Generalization to Broader Scientific Domains
To assess the universality of the ChemBOMAS framework across different scientific domains, we extended our evaluation to a materials science benchmark. This expansion aims to validate a core hypothesis: that the fundamental principle of combining knowledge-driven decomposition with data-driven fine-tuning can be extended beyond chemistry to complex black-box optimization problems. We selected a publicly available dataset representing a unique challenge in scientific discovery:
LNP3 originates from the field of materials science, aiming to address a critical challenge in nanomedicine: optimizing the formulation of lipid nanoparticles (LNPs) for drug delivery. The task involves optimizing LNP composition to effectively encapsulate cannabidiol (CBD). The dataset comprises 768 experimental formulations defined by a 5-dimensional parameter space, encompassing the type and quantity of solid lipids, liquid lipids, and surfactants. This constitutes a complex multi-objective optimization problem requiring the simultaneous achievement of three competing goals: maximizing drug loading and encapsulation efficiency while minimizing particle size. Its parameter space combines categorical and discrete variables, forming a non-trivial search space that represents the challenges encountered in real-world material formulation.
The performance comparison between ChemBOMAS and baseline methods on this dataset is summarized in Table 9.
Dataset Method Best Found Initial Value 95% Max Iter Iteration of Best LNP3 ChemBOMAS 0.62 0.23 12 28 Gollum 0.62 0.21 13 33 BO 0.62 0.25 12 38 LA-MCTS 0.47 0.44 4 15 BO-ICL 0.60 0.15 24 38
In the LNP3 material formulation benchmark, ChemBOMAS demonstrated highly competitive performance. It successfully identified an optimal value of 0.62, matching the final performance achieved by GoLLuM and traditional Bayesian optimization (BO). More importantly, ChemBOMAS demonstrated higher sample efficiency, locating this optimal solution in just 28 iterations, compared to 33 for GoLLuM and 38 for BO. This result indicates that the framework’s structured exploration mechanism can effectively accelerate convergence even in non-chemical optimization scenarios. Notably, while LA-MCTS delivered strong initial performance, it prematurely converged to a suboptimal solution, highlighting the risks of overly aggressive early exploration.
Overall, testing results in the field of materials science demonstrate that ChemBOMAS’s fundamental architecture—namely, the synergistic integration of knowledge-based search space partitioning with data-driven model optimization—holds potential as a universal strategy. It has proven that beyond core chemical domains, this framework possesses equally robust applicability and competitiveness in accelerating black-box optimization across diverse scientific discovery tasks.